### v2

# tidychangepoint: a unified framework for analyzing changepoint detection in univariate time series

Abstract
We present tidychangepoint, a new R package for changepoint detection analysis. Most R packages for segmenting univariate time series focus on providing one or two algorithms for changepoint detection that work with a small set of models and penalized objective functions, and all of them return a custom, nonstandard object type. This makes comparing results across various algorithms, models, and penalized objective functions unnecessarily difficult. tidychangepoint solves this problem by wrapping functions from a variety of existing packages and storing the results in a common S3 class called tidycpt. The package then provides functionality for easily extracting comparable numeric or graphical information from a tidycpt object, all in a tidyverse-compliant framework. tidychangepoint is versatile: it supports both deterministic algorithms like PELT (from changepoint), and also flexible, randomized, genetic algorithms (via GA) that—via new functionality built into tidychangepoint—can be used with any compliant model-fitting function and any penalized objective function. By bringing all of these disparate tools together in a cohesive fashion, tidychangepoint facilitates comparative analysis of changepoint detection algorithms and models.
1 Introduction
1.1 Motivating example
While global temperatures continue to rise, climate scientists continue to face baffling skepticism from some. How do we know that the increase we observe are not just anecdotal? How do we know that the recent pattern of rising temperatures is not simply the latest round of normal fluctuation?
One approach would be to measure temperature over a long period of time, and investigate the time series for the presence of changepoints. Changepoints indicate breaks in a time series that divide it into mutually exclusive regions in which the behavior of the underlying system is the same within each region, but varies across regions. A changepoint might represent a moment at which the system changed in some meaningful and measurable way.
For example, the CET data set collected by Parker, Legg, and Folland (1992) and repackaged in tidychangepoint (Baumer et al. 2024) lists the mean annual temperature in degrees Celsius from 1659 to 2024, as measured in Central England and shown in Figure 1. When did the temperature start to rise? How can we distinguish natural variation in temperature (around a static mean) from a meaningful change in that mean (caused by humans)?
library(tidychangepoint)
plot(CET)
Figure 1: Mean annual temperature in degrees Celsius, as measured in Hadley Centre, England, 1659-2024. Plot created by plot.xts().
Changepoint detection analysis is the study of such changepoints and the algorithms and models that determine them. changepoint (Killick 2024) is a popular R package that implements the PELT algorithm (Killick, Fearnhead, and Eckley 2012) for changepoint detection. We can apply this algorithm to our temperature data using the cpt.meanvar() function.
cet_cpt <- changepoint::cpt.meanvar(as.ts(CET))
In this case, PELT finds 6 changepoints, listed below and illustrated in Figure 2.
index(CET)[cet_cpt@cpts] |>
  format("%Y")
[1] "1713" "1715" "1925" "2002" "2005" "2024"
While PELT has many virtues, there are many alternative changepoint detection algorithms. Another algorithm is Wild Binary Segmentation, provided by the wbs package (Baranowski and Fryzlewicz 2024).
cet_wbs <- wbs::wbs(CET)
index(CET)[sort(cet_wbs\$cpt\$cpt.ic\$mbic.penalty)] |>
  format("%Y")
[1] "1701" "1988"
This algorithms finds 2 changepoints, as seen in Figure 2.
(a) PELT. Plot created by plot.cpt().
(b) WBS. Plot created by plot.wbs().
Figure 2: Changepoints found in temperatue data by two algorithms.
Another promising approach to determining changepoints is to use genetic algorithms that iterate over many possible changepoint sets. Shi et al. (2022) explored the use of a variety of different parametric models using genetic algorithms and a previous version of the CET data set. This analysis leveraged the GA package, which is not specific to changepoint detection, and thus each model that Shi et al. (2022) wanted to test required writing a different set of code, the complexity of which prevents us from reproducing it here. Using a trendshift model (see Section 5), they found the optimal changepoints to be 1700, 1739, and 1988.
There is no optimal solution to the Changepoint Detection Problem (CDP, defined formally in Section 1.2), and in this case, the three algorithms returned different changepoint sets, even when applied to the same data set. Furthermore—and more germaine for us—the objects returned by changepoint detection functions from changepoint, wbs, and GA are all of different types, have difffent structures, contain different information, and can be used with a variety of different functions. This heterogeneity in user interface and package architecture is widespread across roughly a dozen other changepoint detection packages we found. No standardization exists.
These incompatible software interfaces make it unnecessarily difficult to compare the performance of changepoint detection algorithms across different packages. One essentially has to write custom code for each implementation, even when using the same data set. Furthermore, since there is no pre-built interface for genetic changepoint detection algorithms, every tweak to a genetic algorithm for changepoint detection requires custom code.
While we do not claim to offer solutions to the CDP itself, tidychangepoint solves the software heterogeneity problem by providing a unified interface with built-in support for genetic algorithms for changepoint detection. Comparing algorithms and/or models for changepoint detection in tidychangepoint can be as easy as iterating over a list of specifications.
cet_objs <- list(
  null = segment(CET, method = "null"),
  pelt = segment(CET, method = "pelt"),
  wbs = segment(CET, method = "wbs"),
  shi = segment(CET, method = "ga-shi"maxiter = 1000run = 100)
Once the algorithms have been run, we can easily extract the sets of changepoints for comparison using the changepoints() function or compare the results visually using plot() (see Figure 3). Each panel of Figure 3 shows the original time series (black line) segmented by the changepoints determined by the PELT algorithm (vertical dotted lines). The horizontal red lines indicate the mean and 1.96 times the standard deviation within each region, as per the summary produced by tidy() (see Section 3.2).
cet_objs |>
  lapply(plot, use_time_index = TRUEylab = "Degrees Celsius"|>
  patchwork::wrap_plots()
Figure 3: Comparison of four different attempts to detect changepoints in the Central England temperature data. Plots created by plot.tidycpt().
cet_objs |>
  lapply(changepoints, use_labels = TRUE|>
  lapply(as_year)
\$null
character(0)
\$pelt
[1] "1713" "1715" "1925" "2002" "2005"
\$wbs
[1] "1701" "1739" "1740" "1892" "1988"
\$shi
[1] "1691" "1699" "1740" "1741" "1893" "1989"
We return to this example in Section 2.1.
1.2 Mathematical problem formulation
Let $y=\{y_{1},\ldots,y_{n}\}$ be a series of observations made over time points $t=1,\ldots,n$. Define $\tau=\{\tau_{1},\ldots,\tau_{m}\}\subseteq y$ to be a subset of the original observations known as changepoints, for some $0\leq m\leq n$. The $m$ changepoints divide the time series into $m+1$ regions, with the idea that the behavior of the time series within each region does not change, and the changepoints represent points in time at which the behavior of the underlying system changes in some meaningful way. If we adopt the convention that $\tau_{0}=1$ and $\tau_{m+1}=n+1$ then the union of the set of half-open intervals $[\tau_{j},\tau_{j+1})$ for $0\leq j\leq m$ contains the time series $y$.
As there as $2^{n}$ possible changepoint sets, the Changepoint Detection Problem (CDP)—which is the problem of finding the optimal changepoint set (according to some metric)—has attracted considerable attention in the literature. Aminikhanghahi and Cook (2017) provide an excellent survey of relevant results. In unsupervised settings, there is no known true changepoint set, and various models, metrics, and algorithms have been proposed, some of which are in widespread use. We briefly highlight the previous work most relevant to ours in Section 1.3, and review existing software implementations of these algorithms in Section 1.4. In any case, all attempts to solve the CDP considered in this paper involve three essential components: 1) a parametric model $M(y|\theta)$ that describes the data-generating process of the time series; 2) an algorithm $A$ to search the exponentially large space of possible changepoint sets; and 3) a metric $f:(\tau,M)\rightarrow\mathbb{R}$ (a penalized objective function) for evaluating the suitability of any given changepoint set. We define $\tau_{f,M,\theta}^{*}$ to be the changepoint set that minimizes the value of the penalized objective function $f$ on the model $M$ with parameters $\theta$ over the space of all possible changepoint sets. That is,
$\tau_{f,M,\theta_{\tau}}^{*}=\min_{\tau\subseteq y}f(\tau,M(y|\theta_{\tau}))$
Note that the value of the parameters $\theta$ depend on the changepoint set $\tau$, so we may write $\theta_{\tau}$.
For example, perhaps the simplest and most common parametric model for changepoint detection is the normal (Gaussian) meanshift model, in which the time series is assumed to be generated by random variables $Y_{j}\sim N(\mu_{j},\sigma^{2})$ for each region $11\leq j\leq m+$ 1. Thus the series mean $\mu_{j}$ is constant within each of the $m+1$ regions, but varies across the regions, while the variance $\sigma^{2}$ is constant over the entire time series. We fit the normal meanshift model using Maximium Likelihood Estimation (MLE), resulting in the estimated parameters $\hat{\theta}=\{\hat{\sigma}^{2},\hat{\mu}_{0},\ldots,\hat{\mu}_{m}\}$. If we choose the Bayesian Information Criterion (BIC) (Schwarz 1978) as our penalty, then the penalized objective function becomes:
$BIC(\tau,M(y|\hat{\theta}_{\tau}))=k_{M}(\tau)\ln{n}-2\ln{L_{M}(y|\hat{\theta}_{\tau})},$
where $k_{M}(\tau)$ is the number of parameters estimated by the model (in this case, $m+2$) and $L_{M}(y|\hat{\theta}_{\tau})$ is the likelihood function for $M$, which in this case is (see Li and Lund (2012) for details):
$L_{M}(y|\hat{\theta}_{\tau})=-\frac{n}{2}(\ln{\hat{\sigma}^{2}}+1+\ln{2\pi})\,.$
In general, while the true value of $\tau_{BIC,meanshift,\hat{\theta}}^{*}$ is unknown, under reasonable assumptions about the growth of the penalized objective function with respect to the number of changepoints, the PELT algorithm (Killick, Fearnhead, and Eckley 2012) will find the exact value in polynomial time.
More generally, given a time series $y$, a parametric model $M(y|\theta)$, and a penalized objective function $f$, the goal of any changepoint detection algorithm $A$ is to search the exponentially large space of possible changepoints sets for the one $\tau^{*}$ that minimizes the value of $f(\tau,M(y|\hat{\theta}_{\tau}))$. Of course, in the process it must also estimate $\hat{\theta}$.
1.3 Existing changepoint detection algorithms
Aminikhanghahi and Cook (2017) provide a comprehensive survey of different models for changepoint detection in time series. This includes classifications of offline vs. online algorithms, supervised vs. unsupervised problems, computational concerns, robustness, open problems, and the distribution of changepoints within long time series. Guédon (2015) quantifies the uncertainty in the segmentation of diverse sets of changepoints, using both Bayesian and non-Bayesian techniques.
The Pruned Exact Linear Time (PELT) algorithm developed by Killick, Fearnhead, and Eckley (2012) builds on a dynamic programming algorithm from Jackson et al. (2005) to offer a linear time, exact algorithm for changepoint detection. PELT offers order of magnitude improvements in computational cost (i.e., $O(n)$ vs. the $O(n\log{n})$ performance of the binary segmentation algorithm (Scott and Knott 1974) and the $O(n^{3})$ performance of the segmented neighborhood algorithm (Auger and Lawrence 1989)), and is capable of optimizing over different parametric models (e.g., meanvar, meanshift, etc.) and different penalized objective functions (e.g., BIC, mBIC, etc.). To guarantee an optimal solution, PELT requires that the value of the penalized objective function decreases (or increases by an amount bounded by a known constant) with each additional changepoint. Moreover, its running time can be quadratic (i.e., $O(n^{2})$) unless the number of changepoints grows linearly with the length of the time series. While these assumptions are mild, not all possible combinations of time series, models, and penalty functions will meet them.
Li and Lund (2012) illustrate how a genetic algorithm can provide good results by employing Darwinian principles (e.g., survival of the fittest) to search the space of all possible changepoint sets intelligently. For a given parametric model and penalized objective function, a genetic algorithm starts with an initial “generation” of (usually randomly generated) candidate changepoint sets, evaluates them all using the penalized objective function, then probabilistically “mates” the most promising candidate changepoint sets to produce a new generation of the same size. This process is repeated until a stopping criteria is met, and the single best (fittest) changepoint set is returned. While genetic algorithms are neither exact, nor deterministic, nor computationally efficient, they are more easily generalizable than PELT, and can return competitive results in practice (Shi et al. 2022). Li and Lund (2012) also develop the notion of the Minimum Descriptive Length (MDL) as a penalty function based on information theory.
Bai and Perron (1998) use linear regression models fit by least squares to detect changepoints. In particular, they construct a framework for testing the alternative hypothesis that there is one additional changepoint. Barry and Hartigan (1993) employ Bayesian techniques for changepoint detection. Cho and Fryzlewicz (2015) propose the Sparsified Binary Segmentation (SBS) algorithm, which uses CUSUM statistics to analyze high dimensional time series. This algorithm uses the concept of a threshold above which statistics are kept—an idea that resurfaces in Suárez-Sierra, Coen, and Taimal (2023). Cho (2016) explores the utility of a double CUSUM statistic for panel data. Hocking et al. (2013) use machine learning techniques to determine the optimal number of changepoints.
Suárez-Sierra, Coen, and Taimal (2023) detail the implementation of a changepoint detection heuristic we rebrand as Coen’s algorithm in this paper. Coen’s algorithm (Section 4.3) is genetic and uses a non-homogeneous Poisson process model (Section 5.3) to model the exceedances of a threshold over time, and a Bayesian Minimum Descriptive Length (BMDL, Section 6.1.4) penalized objective function. Taimal, Suárez-Sierra, and Rivera (2023) discuss modifications and performance characteristics of Coen’s algorithm.
1.4 Existing changepoint detection R packages
Likely the most prominent R package for changepoint detection is changepoint (Killick 2024; Killick and Eckley 2014), which implements the algorithms PELT, Binary Segmentation, and Segmented Neighborhood. PELT employs one of three different models: meanshift, meanvar, and varshift (see Section 5). For the meanvar model (the most versatile) the data-generating model can follow a Normal, Poisson, Gamma, or Exponential distribution. Penalty functions (see Section 6) implemented by PELT include AIC, BIC, and MBIC, but not MDL. EnvCpt [R-EnvCpt] is another package that wraps changepoint functionality.
The wbs (Baranowski and Fryzlewicz 2024) package implements the Wild Binary Segmentation and standard Binary Segmentation algorithms, the first of which is incorporated into tidychangepoint. The ggchangepoint (Yu 2022) package provides some wrapper functions and graphical tools for changepoint detection analysis that are similar to ours, but more limited in scope and usability. Nonparametric methods for changepoint detection are present in the ecp (James, Zhang, and Matteson 2024) and changepoint.np (Haynes and Killick 2022) packages. We consider these beyond our scope at this time, which is focused on parameteric models for changepoint detection. A Bayesian approach is implemented in the bcp (X. Wang, Erdman, and Emerson 2018; Erdman and Emerson 2008) package, which is no longer available on CRAN. The qcc (Scrucca 2017) package is more broadly focused on quality control and statistical process control, neither of which are specific to changepoint detection. The mcp package focuses on regression-based, Bayesian methods. There are many other changepoint packages, all of which have their strengths and limitations. Please see the “Comparison to other packages” vignette in the mcp documentation for a thorough comparison.
1.5 Our contribution
In this paper, we present several substantive improvements to the existing state-of-the-art. First, the tidychangepoint package (Baumer et al. 2024) provides a unified interface that makes working with existing changepoint detection algorithms in a tidyverse-compliant syntax easy (Section 3). This reduces friction for anyone who wants to compare the results of algorithms, models, or penalized objective functions across different packages. Second, the architecture of the package is easily extended to incorporate new changepoint detection algorithms (Section 4), new parametric models (Section 5), and new penalized objective functions (Section 6). Indeed, our careful separation of these three essential elements of changepoint detection is in and of itself clarifying. Third, to the best of our knowledge, tidychangepoint is the first R package that allows users to run genetic algorithms for changepoint detection without having to write custom code. With GA as a backend, tidychangepoint provides a user-friendly interface for mixing-and-matching various parametric models, penalized objective functions, and initial populations in ways that previously required non-trivial programming by the user. Since tidychangepoint wraps functionality from other packages it adds only modest computational overhead, and delivers correct results (Section 7). Finally, tidychangepoint includes new or updated data sets packaged as time series objects about particulate matter in Bogotá, monthly rainfall in Medellín, temperature in Central England, and batting performance in Major League Baseball. We illustrate the utility of the package in Section 2 and conclude in Section 8.
tidychangepoint is available via download from the Comprehensive R Archive Network (CRAN) since August 19, 2024. The package website contains several short articles demonstrating how to use various features of the package, some of which are not included in this paper (for brevity).
2 Examples using tidychangepoint
In this Section we provide more detailed examples of how tidychangepoint can facilitate meaningful comparisons across algorithms, models, and penalized objective functions.
2.1 Temperatures in Central England
In Section 1.1, we noted that the comparative analysis across different models presented in Shi et al. (2022) required writing custom code for each model fit. In Section 5, we list the many parametric models that tidychangepoint includes. These include the “meanshift” and “trendshift” models, either of which can include AR(1) lagged errors, and either of which can be fit using the BIC or MDL penalties (see Section 6).
With tidychangepoint, we can employ a machine learning style approach to finding the best combination of model and penalty function for a time series. In this case, we build a data.frame with 8 possible combinations.
cet_options <- expand_grid(
  model_fn = c(fit_meanshift_norm, fit_meanshift_norm_ar1,
               fit_trendshift, fit_trendshift_ar1),
  penalty_fn = c(BIC, MDL)
Next, using functionality from the purrr package (Wickham and Henry 2023), we simply iterate the segment() function—which is the main workhorse in tidychangepoint—over these combinations. In each case we employ a genetic algorithm that can produce up to 5000 generations (maxiter), although it will stop if there is no improvement in the fitness after 100 consecutive generations (run).
library(purrr)
cet_mods <- cet_options |>
  mutate(
    tidycpt = map2(
      model_fn, penalty_fn, .f = segment,
      x = CET, method = "ga"maxiter = 5000run = 100
The fitness(), changepoints() and model_name() functions can recover useful information from the resulting objects (see Section 3.2).
cet_mods <- cet_mods |>
  mutate(
    fitness = map_dbl(tidycpt, fitness),
    penalty = map_chr(tidycpt, ~names(fitness(.x))),
    num_cpts = map_int(tidycpt, ~length(changepoints(.x))),
    model_name = map_chr(tidycpt, model_name),,
    num_gens = map_int(tidycpt, ~nrow(as.segmenter(.x)@summary))
We can now easily compare the results and learn about which model performed best with which penalty function.
cet_mods |>
  arrange(penalty, fitness) |>
  select(model_name, fitness, penalty, num_cpts, num_gens)
# A tibble: 8 x 5
  model_name         fitness penalty num_cpts num_gens
  <chr>                <dbl> <chr>      <int>    <int>
1 trendshift            671. BIC            3     1246
2 meanshift_norm_ar1    674. BIC            6     1362
3 meanshift_norm        680. BIC            9     1415
4 trendshift_ar1        682. BIC            4     1347
5 trendshift_ar1        662. MDL            5     2987
6 meanshift_norm        680. MDL            8     1480
7 meanshift_norm_ar1    683. MDL            5     1451
8 trendshift            702. MDL            9     1754
In this case, the trendshift model with white noise errors produces the best fitness score with the BIC penalty, while adding white noise errors produces the best fitness score with the MDL penalty. These results accord with the findings by Shi et al. (2022).
2.2 The designated hitter in Major League Baseball
Sometimes we know the truth about when changepoints occur. For example, Figure 4 shows the difference in batting average between the American and National Leagues of Major League Baseball (MLB) from 1925–2020. The introduction of the designated hitter rule in the American League in 1973 led to noticeable differences in batting performance relative to the National League (which did not adapt the rule until 2020), because the designated hitter employed by American League teams was always a better hitter than the pitcher (who was forced to bat in the National League). Note that in Figure 4, the American League had a higher batting average in every single year from 1973 to 2020. Thus, the year 1973 is a known changepoint, because the data generating process changed before and after that year.
Figure 4: Difference in the rate of home runs per plate appearance between the American and National Leagues in Major League Baseball, 1925–2022. The designated hitter rule was adopted in 1973 by the American League, but not until 2020 by the National League. Note how, during the period from 1973 to 2020 when only the American League employed a designated hitter, the difference in batting average was always negative.
Using the segment() function, we create a list of tidycpt model objects (see Section 3). We apply the PELT algorithm for a change in means and variances using the default MBIC penalty function, and also try a model that allows only the mean to vary across regions and uses the BIC penalty function. Next, we try the WBS algorithm. Finally, we try a genetic algorithm using a normal meanshift model and the MDL penalty function (see Section 6). This algorithm will stop after 500 generations (maxiter), or after 50 consecutive generations with no improvement (run), whichever comes first.
mlb_mods <- list(
  "pelt_meanvar" = segment(mlb_bavg, method = "pelt"),
  "pelt_meanshift" = segment(
    mlb_bavg, method = "pelt",
    model_fn = fit_meanshift_norm, penalty = "BIC"
  "wbs" = segment(mlb_bavg, method = "wbs"),
  "ga" = segment(
    mlb_bavg, method = "ga"model_fn = fit_meanshift_norm,
    penalty_fn = MDL, maxiter = 500run = 50
Because each of the resulting objects are of class tidycpt (see Section 3), comparing the results across the four algorithms is easy using map() (from the purrr package (Wickham and Henry 2023)). First, we examine the changepoint sets.
mlb_mods |>
  map(changepoints, use_labels = TRUE|>
  map(as_year)
\$pelt_meanvar
[1] "1976" "2011"
\$pelt_meanshift
character(0)
\$wbs
[1] "1928" "1930" "1942" "1956" "1972"
\$ga
 [1] "1929" "1931" "1936" "1943" "1948" "1952" "1966" "1969" "1973" "1978"
[11] "2012"
PELT finds two changepoints, including 1976, which is just three years after the true changepoint of 1973. It also finds 2011 to be a changepoint, which is interesting because the difference in batting average climbs noticeably around this time. PELT using the meanshift model and the BIC penalty finds no changepoints. WBS finds 5 changepoints, including 1972, just one year before the true changepoint. Our genetic algorithm finds 11 changepoints, including the true changepoint.
To compare other aspects of the fits, we can iterate the glance() function from the broom package (Robinson, Hayes, and Couch 2024) over that same list. Since glance() always returns a one-row tibble, we can then combine those tibbles into one using the list_rbind() function. Care must be taken to ensure that these comparisons are meaningful. While comparing, say, the elapsed time across these algorithms and models is fair, comparing the fitness values is more fraught. MDL scores are not comparable to BIC or MBIC scores, for example.
mlb_mods |>
  map(glance) |>
  list_rbind()
# A tibble: 4 x 8
  pkg      version algorithm seg_params model_name criteria fitness elapsed_time
  <chr>    <pckg_> <chr>     <list>     <chr>      <chr>      <dbl> <drtn>
1 changep~ 2.3     PELT      <list [1]> meanvar    MBIC       -745.  0.011 secs
2 changep~ 2.3     PELT      <list [1]> meanshift~ BIC        -668.  0.002 secs
3 wbs      1.4.1   Wild Bin~ <list [1]> meanshift~ MBIC          0   0.010 secs
4 GA       3.2.4   Genetic   <list [1]> meanshift~ MDL        -735. 19.982 secs
In Figure 5, we compare the changepoint sets found by the four algorithms. While several of the algorithms indicate a changepoint close to the true changepoint of 1973, it is debatable which provides the best fit to the data.
(a) The PELT algorithm using the mean and variance model and the MBIC penalty.
(b) The PELT algorithm using the meanshift model and the BIC penalty.
(c) The WBS algorithm.
(d) A genetic algorithm using the meanshift model and the MDL penalty.
Figure 5: Comparison of changepoint sets returned by four algorithms. Which provides the best fit to the data?
Visually, we could use a rug plot (see ggplot2::geom_rug()) to show the distribution of changepoints across algorithms (see Figure 6). Three of the four algorithms found a changepoint near the true changepoint of 1973. The other changepoints are false positives with respect to the designated hitter rule, but may very well reflect other changes in batting performance that we simply haven’t been thinking about. We suspect that the recent decrease in the difference in batting average could reflect fewer plate appearances taken by pitchers due to the increased use of relief pitchers.
Figure 6: An update to Figure 4 adding a rug for the distribution of changepoints found by the four algorithms mentioned in this section.
2.2.1 Investigating algorithmic performance
Because tidychangepoint provides model-fitting functions and penalized objective functions that “just work”, it is easy to investigate why candidate changepoint sets of interest may not have been chosen by the algorithm.
Because PELT is known to be optimal under mild conditions, all other candidate changepoint sets should return a value for MBIC that is worse than the one found by PELT (which includes the two changepoints 1976 and 2011). The fitness() function gives us this value.
fitness(mlb_mods[["pelt_meanvar"]])
     MBIC
-744.8353
As expected, the MBIC value of the meanvar model with the true changepoint set is worse (further from $\infty$- ∞).
MBIC(fit_meanvar(mlb_bavg, tau = 49))
[1] -700.8384
Why didn’t PELT pick the changepoint set returned by the genetic algorithm? This set, too, had a worse MBIC score, although its score is better than the true changepoint set.
MBIC(fit_meanvar(mlb_bavg, tau = changepoints(mlb_mods[["ga"]])))
[1] -725.6102
Alternatively, we might ask why the genetic algorithm didn’t pick either the true changepoint set or the one found by PELT. Recall that the genetic algorithm is randomized and only a heuristic, and so unlike in the case of PELT, it is possible that we could manually identify a changepoint set with a lower MDL than the one found by the algorithm. However, in this case, the MDL score for the changepoint set returned by the genetic algorithms bests the other two possibilities.
fitness(mlb_mods[["ga"]])
      MDL
-735.2082
list(49changepoints(mlb_mods[["pelt_meanvar"]])) |>
  map(fit_meanshift_norm, x = mlb_bavg) |>
  map_dbl(MDL)
[1] -691.7307 -682.8757
This type of manual investigation is very difficult to do with other changepoint detection packages.
2.3 Particular matter in Bogotá, Colombia
Consider the time series on particulate matter in Bogotá, Colombia collected daily from 2018–2020, discussed in Suárez-Sierra, Coen, and Taimal (2023), and shown in Figure 7. These data are included in tidychangepoint as bogota_pm.
Figure 7: Particular matter above 2.5 microns in Bogotá, Colombia, recorded daily from 2018–2020.
There appears to be some seasonality in the particulate matter data as well as a pronounced drop in particulate matter around the time of the COVID-19 lockdown in March of 2020. Suárez-Sierra, Coen, and Taimal (2023) use Coen’s algorithm (see Section 4.3) with a manually-specified threshold of 37 $\mu g/m^{3}$ informed by Colombian government policy. Here, we reproduce that result using the GA-based implementation in tidychangepoint. Although we find a different changepoint set, the fit of the NHPP model is of comparable quality.
bog_coen <- bogota_pm |>
  segment(
    method = "ga-coen"maxiter = 1000run = 100,
    model_fn_args = list(threshold = 37)
Alternatively, a careful look at the data suggest that the trendshift model, which allows for a linear slope within each region, might be more appropriate. Setting the method argument of segment() to ga allows us to employ a custom genetic algorithm for which we can specify a model via the model_fn argument. Here, we use the MDL() penalty function and the log_gabin_population() function. While Figure 8 picks up on the COVID-19-related changepoint, Figure 8 shows that the trendshift model also fits the data very well.
bog_ga <- bogota_pm |>
  segment(
    method = "ga"model_fn = fit_trendshift, penalty_fn = MDL,
    population = log_gabin_population(bogota_pm),
    maxiter = 500run = 100
(a) Coen’s algorithm.
(b) A genetic algorithm with a trendshift model.
Figure 8: Changepoint sets for two genetic algorithms applied to the Bogotá particulate matter data. Both pick up on the COVID-19-related changepoint in March of 2020. The latter includes linear trends in each region.
Note also that, without running the genetic algorithm again, we can determine that adding AR(1) lagged errors to our model results in an even lower MDL score.
fitness(bog_ga)
     MDL
8819.181
bogota_pm |>
  fit_trendshift_ar1(tau = changepoints(bog_ga)) |>
  MDL()
[1] 8791.355
The diagnose() function provides an informative visual understanding of the quality of the fit of the model, which may vary depending on the type of model. In Figure 9, note how Figure 9 summarizes the distribution of the residuals, whereas Figure 9 shows the growth of the cumulative number of exceedances (see Section 5.3) relative to the expected growth based on the NHPP model, with the blue lines showing a 95% confidence interval.
(a) The diagnose() function applied to a model object of class nhpp (Coen’s algorithm).
(b) The diagnose() function applied to a model object of class mod_cpt (a genetic algorithm with a trendshift model).
Figure 9: Diagnostic plot for two algorithms applied to the Bogotá particulate matter data. Note how the diagnostic plots differ based on the model type.
3 Architecture of tidychangepoint
The tidychangepoint package implements a simple, flexible structure for working with a myriad of changepoint detection algorithms. tidychangepoint makes extensive use of the S3 object-oriented class system in R. In addition to creating a few new generic functions, it provides methods for a handful of generic functions from other packages (most notably stats and broom) for three new classes of objects (tidycpt, seg_cpt, and mod_cpt), as well as analogous methods for existing objects created by other R packages for changepoint detection (e.g., changepoint). We provide more details in this section.
Every tidycpt object is created by a call to the segment() function. As a verb, “segment” describes the action of splitting a time series into regions, and its usage accords with the tidyverse convention of using verbs as function names. The method argument indicates the algorithm we want to use, and any subsequent arguments (e.g., penalty) are passed to the corresponding function. Here, we pass the penalty argument to force changepoint::cpt.meanvar() to use the BIC penalty function.
mlb_pelt_meanvar_bic <- mlb_bavg |>
  segment(method = "pelt"model_fn = fit_meanvar, penalty = "BIC")
class(mlb_pelt_meanvar_bic)
[1] "tidycpt"
3.1 Object structure
Every tidycpt object contains three sub-objects:
  • segmenter: An object that stores information about an optimal set of changepoints found by an algorithm. A segmenter object could be a cpt object returned by the cpt.meanvar() function from changepoint, a ga object returned by the ga() function from GA, or in principle, any object that provides methods for a few simple generic functions outlined in Section 4.2.
  • model: A model object inheriting from mod_cpt, an internal class for representing model objects. Model objects are created by model-fitting functions, all of whose names start with fit_ (see Section 5). The model of a tidycpt object is the model object returned by the fit_*() function that corresponds to the one used by the segmenter. Because the models described in Section 5 can be fit to any time series based solely on a specified set of changepoints, the information in this object does not depend on the algorithm used to segment the times series: it only depends on the set of changepoints returned by the algorithm.
  • elapsed_time: the clock time that elapsed while the algorithm was running.
3.2 Methods available
tidycpt objects implement methods for the generic functions as.model(), as.segmenter(), as.ts(), changepoints(), diagnose(), fitness(), model_name(), regions, and plot(), as well as three generic functions from the broom package: augment(), tidy(), and glance().
methods(class = "tidycpt")
 [1] as.model     as.segmenter as.ts        augment      changepoints
 [6] diagnose     fitness      glance       model_name   plot
[11] print        regions      tidy
see ’?methods’ for accessing help and source code
For the most part, methods for tidycpt objects typically defer to the methods defined for either the segmenter or the model, depending on the user’s likely intention. To that end, both segmenters and models implement methods for the generic functions as.ts(), changepoints(), model_name(), and nobs().
[1] "as.ts"        "changepoints" "model_name"   "nobs"         "plot"
[6] "print"
The changepoints() function return the indices of the changepoint set, unless the use_labels argument is set to TRUE.
changepoints(mlb_pelt_meanvar_bic)
[1] 15 18 52 87
changepoints(mlb_pelt_meanvar_bic, use_labels = TRUE|>
  as_year()
[1] "1939" "1942" "1976" "2011"
For example, the plot() method for tidycpt simply calls the plot() method for the model of the tidycpt object. The plot shown in Figure 10 uses ggplot2 to illustrate how the proposed changepoint set segments the time series.
plot(mlb_pelt_meanvar_bic)
Figure 10: The MLB time series, as segmented by the PELT algorithm using the normal meanvar model and the BIC penalty.
Changepoint detection models in tidychangepoint follow the design interface of the broom package. Therefore, augment(), tidy(), and glance() methods exist for mod_cpt objects.
  • augment() returns a tsibble (E. Wang, Cook, and Hyndman 2020) that is grouped according to the regions defined by the changepoint set. This object class is appropriate for this use case, since the grouping by changepoint region is natural in this context, and a tsibble is a grouped tibble that is aware of its indexing variable. Users can of course undo the grouping with dplyr::ungroup() or coerce to a regular tibble with tibble::as_tibble().
augment(mlb_pelt_meanvar_bic)
# A tsibble: 95 x 5 [1]
# Groups:    region [5]
   index         y region  .fitted     .resid
   <int>     <dbl> <fct>     <dbl>      <dbl>
 1     1 -0.000173 [1,15) -0.00173  0.00156
 2     2 -0.00166  [1,15) -0.00173  0.0000703
 3     3 -0.00366  [1,15) -0.00173 -0.00192
 4     4  0.000285 [1,15) -0.00173  0.00202
 5     5  0.0105   [1,15) -0.00173  0.0123
 6     6  0.0158   [1,15) -0.00173  0.0175
 7     7 -0.00171  [1,15) -0.00173  0.0000243
 8     8 -0.000439 [1,15) -0.00173  0.00130
 9     9 -0.00650  [1,15) -0.00173 -0.00477
10    10  0.000354 [1,15) -0.00173  0.00209
# i 85 more rows
  • tidy() returns a tibble (Müller and Wickham 2023) that provides summary statistics for each region. These include any parameters that were fit, which are prefixed in the output by param_.
tidy(mlb_pelt_meanvar_bic)
# A tibble: 5 x 10
  region  num_obs      min       max     mean       sd begin   end param_mu
  <chr>     <int>    <dbl>     <dbl>    <dbl>    <dbl> <dbl> <dbl>    <dbl>
1 [1,15)       14 -0.0137   0.0158   -0.00173 0.00779      1    15 -0.00173
2 [15,18)       3 -0.00813 -0.00650  -0.00750 0.000872    15    18 -0.00750
3 [18,52)      34 -0.00923  0.0160    0.00335 0.00628     18    52  0.00335
4 [52,87)      35 -0.0147  -0.000943 -0.00719 0.00315     52    87 -0.00719
5 [87,96)       9 -0.00518 -0.00142  -0.00290 0.00135     87    96 -0.00290
# i 1 more variable: param_sigma_hatsq <dbl>
  • glance() returns a one-row tibble that provides summary statistics for the algorithm. This includes the fitness, which is the value of the penalized objective function (see Section 6) that was used.
glance(mlb_pelt_meanvar_bic)
# A tibble: 1 x 8
  pkg      version algorithm seg_params model_name criteria fitness elapsed_time
  <chr>    <pckg_> <chr>     <list>     <chr>      <chr>      <dbl> <drtn>
1 changep~ 2.3     PELT      <list [1]> meanvar    BIC        -782. 0.009 secs
4 Segmenters
In the example above, the segmenter is of class cpt, because segment() simply wraps the cpt.meanvar() function from the changepoint package.
mlb_pelt_meanvar_bic |>
  as.segmenter() |>
  class()
[1] "cpt"
attr(,"package")
[1] "changepoint"
In addition to the generic functions listed above, segmenters implement methods for the generic functions as.seg_cpt(), fitness(), model_args(), and seg_params().
[1] "as.seg_cpt" "fitness"    "model_args" "seg_params"
Note that while tidychangepoint uses only S3 classes, segmenters (such as those of class cpt) may belong to S4 classes. This is one reason why as.seg_cpt(), which coverts a segmenter of any class to a mod_cpt object, is necessary. fitness() returns a named vector with the type and value of the penalized objective function used by the segmenting algorithm.
fitness(mlb_pelt_meanvar_bic)
      BIC
-782.1165
4.1 Segmenting algorithms provided
Table LABEL:tbl-coverage shows the values passable to the method argument of the segment() function and what they do. Currently, segment() wraps the following algorithms:
  • PELT and Binary Segmentation as implemented by changepoint (Killick 2024)
  • Wild Binary Segmentation as implemented by wbs (Baranowski and Fryzlewicz 2024)
  • A variety of genetic algorithms as implemented by GA (Scrucca 2024). Specific variants include Shi’s algorithm (Shi et al. 2022) and Coen’s algorithm (Suárez-Sierra, Coen, and Taimal 2023).
  • Trivial algorithms for no changepoints, manually input changepoints, and randomly selected changepoints
Table 1: Segmenting methods provided by tidychangepoint.
Method Pkg Segmenter Class Helper Wraps
binseg changepoint cpt NA changepoint::cpt.meanvar()
coen tidychangepoint seg_basket segment_coen() NA
ga GA tidyga segment_ga() GA::ga()
ga-coen GA tidyga segment_ga_coen() segment_ga()
ga-shi GA tidyga segment_ga_shi() segment_ga()
manual tidychangepoint seg_cpt segment_manual() NA
null tidychangepoint seg_cpt segment_manual() NA
pelt changepoint cpt segment_pelt() changepoint::cpt.mean() or changepoint::cpt.meanvar()
random GA tidyga segment_ga_random() segment_ga()
segneigh changepoint cpt NA changepoint::cpt.meanvar()
single-best changepoint cpt NA changepoint::cpt.meanvar()
wbs wbs wbs NA wbs::wbs()
In particular, setting the method argument to the segment() function to ga allows the user to specify any of the model-fitting functions described in Section 5.1 (or a user-defined model-fitting function as in Section 5.2) as well as any of the penalized objective functions in Section 6. While these algorithms can be slow, they are quite flexible.
4.2 Extending tidychangepoint to include new algorithms
In order for a segmenting algorithm to work with tidychangepoint the class of the resulting object must implement, at a minimum, methods for the following generic functions:
  • From stats: as.ts(), and nobs(). These are needed perform basic operations on the time series.
  • From tidychangepoint: changepoints() must return a single integer vector with the best changepoint set as defined by the algorithm. fitness() must return a named double vector with the appropriately named value of the penalized objective function. as.segmenter() dumps the contents of the object into a mod_cpt object. model_name(), model_args(), and seg_params() provide information about the kind of model that was fit, and any arguments passed to the segmenting function.
These compatibility layers are generally not difficult or time-consuming to write.
4.3 Spotlight: Coen’s algorithm
The goal of Coen’s algorithm (Suárez-Sierra, Coen, and Taimal 2023) is to find the optimal set of changepoints for the NHPP model described in Section 5.3. That is, given a time series $y$, find the changepoint set $\tau$ that minimizes the value of $BMDL(\tau,NHPP(y|\hat{\theta}_{\tau}))$. To find good candidates, this genetic algorithm starts with a randomly selected set of popSize changepoint sets, and then iteratively “mates” the best (according to the BMDL penalty function, see Section 6.1.4) pairs of these changepoint sets to recover a new “generation” of popSize changepoint sets. This process is repeated until a stopping criteria is met (most simply, maxiter times). Thus, a constant number (popSize $\times$ maxiter) possible changepoint sets are considered, and the one with the lowest BMDL score is selected.
Coen’s algorithm is implemented in tidychangepoint via the segment() function with the method argument set to ga-coen. As the code below reveals, this is just a special case of the segment_ga() function that wraps GA::ga(). Note that the model_fn argument is set to fit_nhpp and the penalty_fn argument is set to BMDL. The running time of the back-end function GA::ga() is sensitive to the size of the changepoint sets considered, especially in the first generation, and thus we use the population argument to inform the selection of the first generation (see below). Coen’s algorithm runs in about the same time as a naïve algorithm that randomly selects the same number of changepoints, but produces changepoint sets with significantly better BMDL scores.
segment_ga_coen
function(x, ...) {
  segment_ga(
    x, model_fn = fit_nhpp, penalty_fn = BMDL,
    population = build_gabin_population(x), popSize = 50, ...
<bytecode: 0x62d139cfb7f0>
<environment: namespace:tidychangepoint>
By default, the function GA::gabin_Population() selects candidate changepoints uniformly at random with probability 0.5, leading to candidate changepoint sets of size $n/2$, on average. These candidate changepoint sets are usually poor (since $n/2$ changepoints is ludicrous), and we observe much better and faster performance by seeding the first generation with smaller candidate changepoint sets. To this end, the build_gabin_population() function runs several fast algorithms (i.e., PELT, Wild Binary Segmentation, etc.) and sets the initial probability of being selected to three times the average size of the changepoint set found by these algorithms. This results in much more rapid covergence around the optimal changepoint set. Alternatively, tidychangepoint also provides the log_gabin_population() function, which sets the initial probability to $\ln{n}/n$ (see Section 7.2).
5 Models
All model objects are created by calls to one of the model-fitting functions listed in Section 5.1, whose name begins with fit_. These functions all inherit from the class fun_cpt and can be listed using ls_models(). All model objects inherit from the mod_cpt base class.
The model object in our example is created by fit_meanvar(), and is of class mod_cpt.
mlb_pelt_meanvar_bic |>
  as.model() |>
  str()
List of 6
 \$ data         : Time-Series [1:95] from 1 to 95: -0.000173 -0.001663 -0.003656 0.000285 0.010521 ...
 \$ tau          : int [1:4] 15 18 52 87
 \$ region_params: tibble [5 x 3] (S3: tbl_df/tbl/data.frame)
  ..\$ region           : chr [1:5] "[1,15)" "[15,18)" "[18,52)" "[52,87)" ...
  ..\$ param_mu         : num [1:5] -0.00173 -0.0075 0.00335 -0.00719 -0.0029
  ..\$ param_sigma_hatsq: Named num [1:5] 5.63e-05 5.07e-07 3.83e-05 9.67e-06 1.62e-06
  .. ..- attr(*, "names")= chr [1:5] "[1,15)" "[15,18)" "[18,52)" "[52,87)" ...
 \$ model_params : NULL
 \$ fitted_values: num [1:95] -0.00173 -0.00173 -0.00173 -0.00173 -0.00173 ...
 \$ model_name   : chr "meanvar"
 - attr(*, "class")= chr "mod_cpt"
In addition to the generic functions listed above, models implement methods for the generic functions coef(), fitted(), logLik(), plot(), and residuals(), as well as augment(), tidy(), and glance().
[1] "augment"   "coef"      "diagnose"  "fitted"    "glance"    "logLik"
[7] "regions"   "residuals" "tidy"
Like other model objects that implement glance() methods, the output from the glance() function summarizes the fit of the model to the data.
mlb_pelt_meanvar_bic |>
  as.model() |>
  glance()
# A tibble: 1 x 11
  pkg   version algorithm params num_cpts    rmse logLik   AIC   BIC  MBIC   MDL
  <chr> <pckg_> <chr>     <list>    <int>   <dbl>  <dbl> <dbl> <dbl> <dbl> <dbl>
1 tidy~ 1.0.0.~ meanvar   <NULL>        4 0.00507   367. -706. -671. -689. -674.
5.1 Model-fitting functions provided
Table LABEL:tbl-models lists the model-fitting functions provided by tidychangepoint, along with their parameters. Any of these functions can be passed to the model_fn argument of segment_ga() (see Section 4.1).
Table 2: Model-fitting functions provided by tidychangepoint. Unless otherwise noted, all models assume a normal distribution and white noise errors. The suffix _ar1 indicates autocorrelated AR(1) lagged errors (as decribed in Shi et al. (2022)), where $\phi$ is the measure of autocorrelation. See Section 5.3 for a description of the non-homogeneous Poisson process model used by fit_nhpp().
Function name $k$ Parameters ($\theta$) Note
fit_meanshift_norm() $m+2$ $\sigma^{2},\mu_{0},\ldots,\mu_{m}$ Normal
fit_meanshift_lnorm() $m+2$ $\sigma^{2},\mu_{0},\ldots,\mu_{m}$ log-normal
fit_meanshift_pois() $m+2$ $\sigma^{2},\mu_{0},\ldots,\mu_{m}$ Poisson
fit_meanshift_norm_ar1() $m+3$ $\sigma^{2},\phi,\mu_{0},\ldots,\mu_{m}$
fit_trendshift() $32m+$ 3 $\sigma^{2},\mu_{0},\ldots,\mu_{m},\beta_{0},\ldots,\beta_{m}$ special case of fit_lmshift() with $p=1$
fit_trendshift_ar1() $42m+$ 4 $\sigma^{2},\phi,\mu_{0},\ldots,\mu_{m},\beta_{0},\ldots,\beta_{m}$
fit_lmshift() $p(m+1)+1$ $\sigma^{2},\beta_{00},\ldots,\beta_{0m},\beta_{10},\ldots,\beta_{pm}$ polynomial [^1] Unlike fit_meanshift_norm(), fit_lmshift() function uses the lm() function from the stats, which provides flexibility at the expense of speed (fit_meanshift_norm() is faster). of degree $p$
fit_lmshift_ar1() $p(m+1)+1$ $\sigma^{2},\beta_{00},\ldots,\beta_{0m},\beta_{10},\ldots,\beta_{pm}$
fit_meanvar() $22m+$ 2 $\sigma_{0}^{2},\ldots,\sigma_{m}^{2},\mu_{0},\ldots,\mu_{m}$
fit_nhpp() $22m+$ 2 $\alpha_{0},\ldots,\alpha_{m},\beta_{0},\ldots,\beta_{m}$
5.2 Extending tidychangepoint to include new models
Users can create their own model-fitting functions. The names of these functions should start with fit_ for consistency, and they must be registered with a call to fun_cpt(). The first argument x must be the time series and the second argument tau must be a subset of the indices of x. Any subsequent arguments can be passed through the dots. Every fit_ function returns an object of class mod_cpt, which can be created with the arguments shown below.
args(new_mod_cpt)
function (x = numeric(), tau = integer(), region_params = tibble::tibble(),
    model_params = double(), fitted_values = double(), model_name = character(),
NULL
5.3 Spotlight: Non-homogenous Poisson process model
Having already described the meanshift model in Section 1, we now describe mathematically the NHPP model used in Coen’s algorithm (see Section 4.3) and created by the fit_nhpp() function listed in Section 5.1.
Let $z_{w}(y)=\{t:y_{t}>w\}$ be a subset of the indices of the original times series $y$, for which the observation $y_{t}$ exceeds some threshold $w$, which by default is the empirical mean $\bar{y}$. This new time series $z_{w}(y)$ is called the exceedances of the original time series $y$. Following Suárez-Sierra, Coen, and Taimal (2023), we model the time series $z_{w}(y)$ as a non-homogenous Poisson process (NHPP), using a Weibull distribution parameterized by shape ($\alpha$) and scale ($\beta$). In a Bayesian setting, those parameters follow a Gamma distribution with their appropriate hyperparameters. Thus, we assume that there exists a set of parameters $\theta=(\alpha_{0}\ldots\alpha_{m},\beta_{0},\ldots,\beta_{m})$ for which the exceedances $z_{w}(y)$ of the original time series $y$ are modeled accurately as a non-homogenous Poisson process.
The best-fit parameters are determined by maximizing the log posterior function described below:
$\ln{g(\theta_{\tau}|y)}\propto\ln{L_{NHPP}(y|\theta_{\tau})}+\ln{g(\theta_{\tau})}\,.$
where $L$ is the likelihood function and $g$ is the prior probability distribution function. Closed-form expressions for each of these components for a Weibull distribution with two Gamma priors are given in Suárez-Sierra, Coen, and Taimal (2023). The tidy() function displays the values of the fitted parameter values, or one can pull the region_params object directly from the model object.
bog_coen |>
  as.model() |>
  pluck("region_params")
# A tibble: 4 x 5
  region             param_alpha param_beta logPost logLik
  <chr>                    <dbl>      <dbl>   <dbl>  <dbl>
1 [1,379)                  0.711     0.808   -202.  -198.
2 [379,820)                0.661     0.0762  -346.  -344.
3 [820,1.03e+03)           0.484     0.0831   -42.6  -40.6
4 [1.03e+03,1.1e+03)       0.668     0.0788   -48.4  -46.6
6 Penalized objective functions
We call the function $f$ a penalized objective function [^2] Occasionally, we may abuse terminology by referring to $f$ as a penalty function, but technically, $P_{f}$ is the penalty function. because it comes in the form
$f(\tau,M(y|\theta_{\tau}))=P_{f}(\tau,n)-2\ln{L_{M}(y|\theta_{\tau})}\,,$
where $P_{f}$ is a function that guards against overfitting by penalizing large changepoint sets—that is, it is a monotonically increasing function of $m$. The stats package (R Core Team 2024) provides generic methods for the well-known penalty functions AIC() and BIC(), which work seamlessly on tidychangepoint models because they all implement methods for logLik(). In addition, tidychangepoint provides generic functions and methods for four additional penalty functions: HQC(), MBIC(), MDL(), and BMDL(). For compatibility with changepoint we also provide the alias SIC() for BIC(). tidychangepoint supports all of the penalty functions supported by changepoint, with the exception of the “Asymptotic” and “CROPS” penalties, which are not fully generalizable.
All penalty functions return 0 for the special case in which $m=0$. For ease of explanation, let the vector $\ell_{j}=\tau_{j+1}-\tau_{j}$ for $11\leq j\leq m+$ 1 encode the lengths of the $m+1$ regions defined by the changepoint set $\tau$.
6.1 Penalized objective functions supported in tidychangepoint
6.1.1 Hannan-Quinn Information Criterion
The Hannan-Quinn information Criterion is a slight tweak on the BIC that involves an iterated logarithm (Hannan and Quinn 1979): $P_{HQC}(\tau,n)=2m\cdot\ln{\ln{n}}$.
6.1.2 Modified Bayesian Information Criterion
Following Zhang and Siegmund (2007) and Li and Lund (2012), we define the MBIC as:
$P_{MBIC}(\tau,n)=3m\ln{n}+\sum_{j=1}^{m+1}\ln{\frac{\ell_{j}}{n}}\,.$
6.1.3 Minimum Descriptive Length
As described in Shi et al. (2022) and Li and Lund (2012), we define the MDL as:
$P_{MDL}(\tau,n)=\frac{a(\theta_{\tau})}{2}\cdot\sum_{j=0}^{m}\log{\ell_{j}}+2\ln{m}+\sum_{j=2}^{m}\ln{\tau_{j}}+\left(2+b(\theta_{\tau})\right)\ln{n}\,.$
where $a(\theta)$ is the number of parameters in $\theta$ that are fit in each region, and $b(\theta)$ is the number of parameters fit to the model as a whole. For example, the in the meanshift model, $a(\theta)=1$ to account for the $\mu_{i}$’s, and $b(\theta)=1$ to account for $\sigma^{2}$.
The MBIC and MDL differ from the penalties imposed by either the Akaike Information Criterion (AIC) or the BIC in important ways. For example, the penalty for the AIC depends just linearly on the number of changepoints, while the BIC depends on the number of changepoints as well as the length of the original times series. Conversely, the MDL penalty depends not only on the number of changepoints, but the values of the changepoints. Because of this, changepoints far from zero can have disproportionate impact.
6.1.4 Bayesian Minimum Descriptive Length
The Bayesian Minimum Descriptive Length combines the MDL penalty function with the log prior $g$ for the best-fit parameters $\hat{\theta}$ in the NHPP model described in Section 5.3. Currently, the BMDL penalty can only be applied to the NHPP model.
The exact value of the BMDL is then:
$BMDL(\tau,NHPP(y|\hat{\theta}_{\tau}))=P_{MDL}(\tau)-2\ln{L_{NHPP}(y|\hat{\theta}_{\tau})}-2\ln{g(\hat{\theta}_{\tau})}\,.$
6.2 Extending tidychangepoint to include new penalty functions
New penalty functions can be contributed to tidychangepoint by defining a new generic function and implementing a method for the logLik class. Penalty functions should return a double vector of length one.
Note that similar to the logLik() method for lm, the logLik() method for mod_cpt embeds relevant information into the object’s attributes. These include the quantities necessary to compute AIC(), BIC(), MDL(), etc.
mlb_pelt_meanvar_bic |>
  as.model() |>
  logLik() |>
  unclass() |>
  str()
 num 367
 - attr(*, "num_params_per_region")= int 2
 - attr(*, "num_model_params")= int 0
 - attr(*, "df")= num 14
 - attr(*, "nobs")= int 95
 - attr(*, "tau")= int [1:4] 15 18 52 87
7 Results
In this section we document tidychangepoint’s correctness and computational costs.
7.1 Correctness
Since tidychangepoint relies on other packages to actually segment time series via their respective changepoint detection algorithms, the results returned by segment() are correct to the extent that the dependent results are correct.
However, recall that the model component of a tidycpt object is computed by our fit_*() model-fitting functions, after the segmenting algorithm has returned its results. We show below that these computations are accurate.
First, we match the model parameters returned by fit_trendshift() and fit_trendshift_ar1() with the results for the Central England Temperature data (CET) reported in rows 1–4 of Table 2 in Shi et al. (2022). Unfortunately, the CET data been revised since the publication of Shi et al. (2022), so the temperatures present in CET do not exactly match the temperatures used in Shi et al. (2022) for all years. Accordingly, the numbers in Table LABEL:tbl-shi-compare do not match exactly, but they are very close. However, when we perform this same analysis on the old data (not shown here), the numbers match exactly. [^3] Other results from Table 3 also match.
cpts <- c(170017391988)
ids <- time2tau(cpts, as_year(time(CET)))
trend_wn <- CET["/2020-01-01"|>
  fit_trendshift(tau = ids)
trend_ar1 <- CET["/2020-01-01"|>
  fit_trendshift_ar1(tau = ids)
Table 3: Comparison of results reported by Shi et al. (2022) to those reported by tidychangepoint. All discrepancies are caused by revisions to the underlying CET data used.
Model $\hat{\sigma}^{2}$ logLik BIC MDL $\hat{\phi}$
Shi et al. (2022) WN 0.291 -290.02 650.74 653.07 NA
trend_wn 0.290 -289.86 650.41 652.75 NA
Shi et al. (2022) AR(1) 0.290 -288.80 654.19 656.52 0.058
trend_ar1 0.290 -288.50 653.59 655.94 0.055
Second, in Table LABEL:tbl-param-compare we compare the estimated parameter values from the model component of a tidycpt object (which are fit by fit_meanvar()) to those from the segmenter component (which are fit by cpt.meanvar()) using a simulated data set.
tidycpt <- DataCPSim |>
  segment(method = "pelt"penalty = "BIC")
# Results shown in Table...
# tidycpt\$segmenter@param.est
# y\$model\$region_params
Table 4: Comparison of parameter values returned by fit_meanvar() ($\hat{\mu}_{tidycpt}$ and $\hat{\sigma}^{2}_{tidycpt}$) to those of changepoint::cpt.meanvar() ($\hat{\mu}_{cpt}$ and $\hat{\sigma}^{2}_{cpt}$). The discrepancies in the means are caused by tidychangepoint using intervals open on the right, as opposed to the left. Additional discrepancies in the variances are caused by changepoint’s use of the population variance.
region $\hat{\mu}_{tidycpt}$ $\hat{\sigma}^{2}_{tidycpt}$ $\hat{\mu}_{cpt}$ $\hat{\sigma}^{2}_{cpt}$
[1,547) 35.28 127.09 35.28 126.88
[547,822) 58.10 371.75 58.20 370.52
[822,972) 96.71 923.95 96.77 920.98
[972,1.1e+03) 155.85 2441.82 156.52 2405.97
The tiny differences in the means are because tidychangepoint uses intervals that are closed on the left and open on the right, whereas changepoint does the opposite. The differences in the variances are caused by the same issue and changepoint’s reporting of the population variance instead of the sample variance.
7.2 Benchmarking
Since segment() wraps underlying functionality from other packages, there is overhead created, both in terms of memory and speed. Table LABEL:tbl-bench shows the results of benchmarking segment() against changepoint::cpt.meanvar(), which is doing the underlying work in both cases. In this case, segment() uses about twice as much memory and is about 8 times slower than the underlying function. While that may seem like a lot, we’re talking about microseconds.
Table 5: Benchmarking
expression min median itr/sec mem_alloc gc/sec n_itr n_gc total_time
tidychangepoint 13.05ms 13.99ms 71.59874 75.7KB 0.0000 10 0 139.7ms
changepoint 1.45ms 1.56ms 595.09531 35.7KB 66.1217 9 1 15.1ms
As noted in Section 4.3, tidychangepoint provides two alternative functions to GA::gabin_Population() for creating an initial population for a genetic algorithm: build_gabin_population() and log_gabin_population(). Both perform some näive analysis to set the probability of an observation being selected uniformly at random to belong to the first generation. The former runs three quick algorithms (PELT, Binary Segmentation, and WBS) and sets the probability to three times the average size of the changepoint sets returned by those algorithms, divided by $n$. The latter sets the probability to $\log{n}/n$. Both functions can be thought of as “informed” by the data, albeit in different ways. Recall that GA::gabin_Population() always uses the probability $1/2$, and can thus be considered “uninformed”.
In Figure 11, we compare the performance of a genetic algorithm on monthly retail data from Australia (O’Hara-Wild et al. 2022), using these three methods of forming the initial population. In this test, the maximum number of iterations was set to 1000, with a consecutive run of 100 generations with no improvement stopping the algorithm. We see in Figure 11 that both of the “informed” initial populations result in improved performance. The recovered changepoint sets score better values of the penalized objective function, and both algorithms terminated well before they reached the maximum number of generations. Conversely, the algorithm using the default initial population was still going after 1000 generations. In fact, it only took 35 generations for the BIC scores returned by the genetic algorithm whose initial population was set by log_gabin_population() to beat the best score returned by the same algorithm using the default initial population after 1000 generations.
Figure 11: Comparison the evolution of the best candidate changepoint sets for genetic algorithms with different initial populations. The ‘runif’ series begins by selecting observations to belong to the candidate changepoint set uniformly at random, with probability 0.5. On the other hand, the ‘build’ and ‘logn’ series also selects candidate changepoints uniformly at random, but with informed probabilities.
Of course, it is possible that by casting a wider net, the default initial population will catch changepoint sets that were randomly excluded by the “informed” initial populations. But these results held up under repeated testing, and they suggest a promising line of attack for future work.
8 Conclusion
8.1 Future work
There is more work to be done to bring additional changepoint detection algorithms, models, and penalty functions into tidychangepoint. Please see the package’s GitHub Issues page for a complete list.
Generally, computational costs for changepoint detection algorithms are a concern. Some algorithms have a known theoretical running time in relation to the length of the time series, while others do not. Moreover, while the implementations of some algorithms (notably, changepoint and GA) use pre-compiled C code, others may not. All of the code for tidychangepoint is written in R, but since it is mainly an interface that wraps code from other packages, the overhead is slight (see Section 7.2), and there isn’t much that can be done to improve performance. Even the new algorithms that are provided by tidychangepoint (e.g., Coen’s algorithm, Shi’s algorithm) are simply special cases of genetic algorithms that use GA as a backend.
Another major issue with changepoint detection algorithms is the robustness of the results to a variety of changes in conditions, assumptions, or computational environments. While some algorithms report the probability of an observation being a changepoint, others do not. Some models rely on assumptions about the distribution or frequency of changepoints relative to the length of the time series. In randomized algorithms (such as genetic algorithms), the robustness of the results may be interrelated with stopping criteria, the initial seed or population, or other computational considerations. CUSUM methods operate on a summary of the original time series: are they less susceptible to outliers? These issues are not correctable within tidychangepoint. We will continue to engage with the authors of the underlying packages when we see areas for improvement and alert our own users to inconsistencies through our documentation.
Finally, in this paper we have considered the multiple changepoint detection problem for a univariate time series, but multivariate time series are also being studied in a variety of contexts.
8.2 Summary
We have described how tidychangepoint offers a flexible but standardized architecture for changepoint detection analysis in R in a manner that is consistent with the design principles of the tidyverse. Previously, changepoint detection analysis could be conducted using a variety of R packages, each of which employed its own non-conforming interface. Our unified interface allows users to compare results across different changepoint algorithms, models, and penalty functions with ease.